OpenStudy (anonymous):

What is the area of a triangle with a base of 6 units and a height of 14 units? Ig necessary, round your answer to two decimal places.

OpenStudy (anonymous):

Area of a triangle : \[\frac{1}{2} * Base * height\]

OpenStudy (anonymous):

Substitute the values and you'd get the answer ! :)

OpenStudy (anonymous):

I got 42??? is that right?

OpenStudy (anonymous):

3*14=42 you still have to multiply by 1/2 42*1/2= ?

OpenStudy (anonymous):

21?

OpenStudy (anonymous):

yupp!

OpenStudy (anonymous):

Thankk youu